OpenStudy (luigi0210):

I just need someone to check me and see if I'm right so far.. Find the centroid of the region bounded by: y=sinx, y=cosx, x=0, \(x=\frac{\pi}{4}\)

OpenStudy (luigi0210):

*\[ (x_{c},~y_{c}) =(\frac{\int(x)(cosx-sinx)dx}{\int (cosx-sinx)dx}, ~\frac{\frac{1}{2}\int(cosx+sinx)(cosx-sinx)dx}{\int(cosx-sinx)dx})\]

OpenStudy (anonymous):

NO NO NO NO...

OpenStudy (luigi0210):

Where did I screw up? >_>